# Evaluating GPT-V4 (GPT-4 with Vision) on Detection of Radiological Findings on Chest Radiographs

### v1

@@ Results
In the zero-shot setting, in the task of detecting ICD-10 codes alone, GPT-4V attained a G&R+/R+ of 12.3%, a G&R+/G+ of 5.8%, and an F1 score of 7.3% on the NIH dataset, and a G&R+/R+ of 25.0%, a G&R+/G+ of 16.8%, and an F1 score of 18.2% on the MIDRC dataset. When both the ICD-10 codes and their corresponding laterality were considered, GPT-4V produced a G&R+/R+ of 7.8%, a G&R+/G+ of 3.5%, and an F1 score of 4.5% on the NIH dataset, and a G&R+/R+ of 3.6%, a G&R+/G+ of 4.9%, and an F1 of 6.4% on the MIDRC dataset. In few-shot learning, GPT-4V showed improved performance on both datasets. When contrasting zero-shot and few-shot learning, there were improved G&R+/G+ and F1 scores in few-shot learning. Nonetheless, there was not a substantial increase in the G&R+/R+.
@@ 2.3 Experimental setup
To obtain the reference standard tables, GPT-4 was used to convert each free-text radiology report (Figure 2a) into a table of radiological findings with the prompt shown in Figure 2b. This table included the radiological findings, the corresponding ICD-10 diagnostic codes and their laterality, and detailed descriptions of the ICD codes (Figure 2c). Subsequently, each report was independently evaluated by three readers from a cohort of five board-certified radiologists and residents (H.O., P.K., C.W., J.K., G.S). Two of the readers are third-year radiology residents, and the remaining three are radiology attendings, each with over 15 years of experience. Their radiology subspecialties cover chest, ED/bone, neurology, and body imaging. All readers had access to the image views and reports but not to additional clinical or patient data. Both datasets, the National Institutes of Health (NIH) and the Medical Imaging and Data Resource Center (MIDRC), were comprehensively reviewed, with each dataset’s 50 reports being examined by three reviewers to maintain consistency and objectivity in the evaluation process. Findings were only included in the final tables if they were observed by at least two of the three reviewers; conversely, findings were excluded if two or more reviewers did not identify them. The majority vote principle was employed to provide a clear consensus for the presence or absence of radiological findings and lead to the final reference-standard table.
@@ 2.3.1 Evaluating Performance in the zero-shot setting
To evaluate GPT-4V’s performance in the zero-shot setting, the chest radiographs (Figure 2d) were inputted into the GPT-4V model (version from Oct.13th)[2] along with the prompt (Figure 2e). This step aims to generate a radiological findings table (Figure 2f) comparable to the reference standard. The analysis concentrated on aligning positive radiological finding identification and laterality between GPT-generated and consensus tables. G&R+/R+ (positive predictive value), G&R+/G+ (true positive rate), and F1 scores were calculated at the report level (detailed in the section of “Evaluation metrics”). These metrics were used to assess the accuracy of GPT-4V in detecting the ICD-10 codes and their respective laterality.
@@ 2.3.2 Evaluating Performance in the few-shot setting
To evaluate GPT-4V’s performance in the few-shot setting, the input was extended to include two examples of chest radiographs with their corresponding radiological findings tables before the prompt. From the pool of 10 chest radiographs, two were randomly selected to serve as few-shot examples for the GPT-4V model. Supplying the model with these examples helped to provide context, boosting the model’s capacity to generate an accurate radiological findings table. The same performance metrics— G&R+/R+, G&R+/G+, and F1 score— were used at the report level to assess the effectiveness of few-shot learning.
@@ 2.4 Statistical Analysis
First, the GPT-4V was employed to detect radiological findings from each chest radiography and compared the predicted findings with those obtained from the radiologists. To convey the performance evaluation, G&R+ (true positive rate) was used to represent the number of ICD-10 Codes correctly predicted by GPT-4V, R+ for the total number of ICD-10 Codes identified by GPT-4V, and G+ for the count of ICD-10 Codes existing in the reference standard.
Evaluation metrics of G&R+/R+ (positive predictive value, PPV), G&R+/G+ (true positive rate, TPR), and F1 score were used to assess GPT-4V’s performance in detecting imaging findings from each chest radiograph [6]. These predicted findings were then compared with those obtained from the radiologists.
@@
For example, in Figure 2, while evaluating ICD-10 codes alone, there were 2 G&R+ (J90 and J18.9), five findings predicted by GPT-4V, and four in the reference standard table. Therefore, G&R+/R+ was 0.4 (=2/(2+3)) , G&R+/G+ was 0.5 (=2/(2+2)), and F1 score was 0.44.
After obtaining the G&R+/R+, G&R+/G+, and F1 score for each chest radiography, the macro averages were calculated for G&R+/R+, G&R+/G+, and F1 score across all 90 chest radiographs.
@@ 3.1 Performance in the zero-shot setting
The performance of the GPT-4V model in the zero-shot setting varied across the NIH and MIDRC datasets and under different circumstances (Figure 3). In the task of detecting ICD-10 codes alone, the model attained a G&R+/R+ of 5.53/45 (12.3%), with a standard deviation (SD) of 0.25, a standard error (SE) of 0.04, and an interquartile range (IQR) of 0.20; a G&R+/G+ of 2.60/45 (5.8%), with a SD of 0.10, a SE of 0.02, and an IQR of 0.10; and an F1 score of 3.30/45 (7.3%), with a SD of 0.13, a SE of 0.02, and an IQR of 0.14, on the NIH dataset. Conversely, on the MIDRC dataset, the model managed a G&R+/R+ of 11.25/45 (25.0%), with a SD of 0.24, a SE of 0.04, and an IQR of 0.33; a G&R+/G+ of 7.56/45 (16.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.25; and an F1 score of 8.20/45 (18.2%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.29. The notable differences in measurements between the two datasets were primarily due to the MIDRC dataset having fewer missing GPT-4V-generated ICD-10 codes than the NIH dataset. On the MIDRC dataset, GPT-4V generated 144 radiological findings, while the reference standard comprised 261 findings. However, on the NIH dataset, GPT-4V produced 102 radiological findings, while the reference standard comprised 220 findings. Nevertheless, when both the ICD-10 codes and their corresponding laterality were taken into account, the GPT-4V model in the zero-shot setting produced a G&R+/R+ of 3.5/45 (7.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.0; a G&R+/G+ of 1.56/45 (3.5%), with a SD of 0.01, a SE of 0.01, and an IQR of 0.0; and an F1 score of 2.05/45 (4.5%) with a SD of 0.10, a SE of 0.02, and an IQR of 0.0 on the NIH dataset; and a G&R+/R+ of 4.92/45 (3.6%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.25; a G&R+/G+ of 2.19/45 (4.9%), with a SD of 0.09, a SE of 0.01, and an IQR of 0.10; and an F1 of 2.90/45 (6.4%), with a SD of 0.11, a SE of 0.02, and an IQR of 0.15, on the MIDRC dataset.
Figure 3: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the zero-shot setting, with statistical significance assessed using a two-tailed t-test. a, Detecting the radiological findings in ICD codes. b, Detecting both the radiological findings in ICD codes and their corresponding lateralities. * - p < 0.05; ** - p < 0.01; *** - p < 0.001; ns – Not significant. ICD= International Classification of Diseases, Tenth Revision.
@@ 3.2 Performance in the few-shot setting
In few-shot learning, GPT-4V showed improved performance on both NIH and MIDRC datasets (Figure 4). When the model was provided with two illustrative chest radiographs and reports, there was a marked improvement in the G&R+/R+ on the NIH dataset, with rates rising to 5.72/45 (12.7%) with a SD of 0.19, a SE of 0.03, and an IQR of 0.25. G&R+/G+ also enhanced to 4.69/45 (10.4%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22; and the F1 score reached 5.03/45 (11.1%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22. On the MIDRC dataset, the G&R+/R+ was improved to 16.15/45 (35.9%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; the G&R+/G+ improved to 16.68/45 (37.1%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; and the F1 score improved to 15.47/45 (34.3%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50. When tasked with detecting both ICD-10 codes and their corresponding lateralities, the GPT-4V model demonstrated improved efficacy. On the NIH dataset, it displayed a G&R+/R+ of 1.62/45(3.5%), with a SD of 0.10, a SE of 0.01, and an IQR of 0.0; a G&R+/G+ of 1.14/45 (2.5%), with a SD of 0.07, a SE of 0.01, and an IQR of 0.0; and an F1 score of 1.30/45(2.8%), with a SD of 0.08, a SE of 0.01, and an IQR of 0.0. On the MIDRC dataset, it achieved a G&R+/R+ of 8.96/45 (19.9%), with a SD of 0.22, a SE of 0.03, and an IQR of 0.33; a G&R+/G+ of 9.14/45 (20.3%), with a SD of 0.25, a SE of 0.04, and an IQR of 0.33; and an F1 score of 8.53/45 (19.0%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.31.
Figure 4: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the few-shot setting, with statistical significance assessed using two-tailed t-test. a, Detecting the radiological findings in ICD codes. b, Detecting both the radiological findings in ICD codes and their corresponding lateralities. * - p < 0.05; ** - p < 0.01; *** - p < 0.001; **** - p < 0.0001; ns – Not significant. ICD= International Classification of Diseases, Tenth Revision.
When contrasting zero-shot and few-shot learning approaches on both datasets (Figure 5), there were improved G&R+/G+ and F1 scores in few-shot learning in both scenarios (ICD-10 codes only and ICD-10 codes with laterality). Nonetheless, there was not a substantial increase in the G&R+/R+ suggesting that while few-shot learning may enhance the model’s capacity to detect findings, it does not noticeably enhance the precision of the predictions.
@@ 4 Discussion
The emergence of multimodal large language models (LLMs) that can understand both text and images, such as GPT-4 with Vision (GPT-4V), shows potential for automated image-text pair generation. However, applying these models to real-world data is yet to be thoroughly examined. This study assessed the feasibility of using GPT-4V to detect radiological findings from chest radiographs in both zero-shot and few-shot learning contexts. The results (G&R+/R+ of 5.53/45 (12.3%) and G&R+/G+ of 2.60/45 (5.8%)) demonstrated that radiological findings tables generated by the GPT-4V model still need to be prepared for use in clinical practice. We acknowledged a limitation in employing GPT-4 for converting radiological reports into a structured table, where inaccuracies in ICD-10 code assignments and distinguishing between radiological findings and conclusions may impact the data’s reliability and interpretability. A notable limitation of the GPT-4V output was its failure to detect several radiological conditions based on corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10). Overall, the top three findings that GPT-4V could not detect were “endotracheal tube”, “central venous catheter”, and “degenerative changes of osseous structures”. Conversely, GPT-4V most accurately detected findings such as “chest drain”, “air-space disease”, and “lung opacity”.

### v2

@@ Results
In the zero-shot setting, in the task of detecting ICD-10 codes alone, GPT-4V attained an average PPV of 12.3%, an average TPR of 5.8%, and an average F1 score of 7.3% on the NIH dataset, and an average PPV of 25.0%, an average TPR of 16.8%, and an average F1 score of 18.2% on the MIDRC dataset. When both the ICD-10 codes and their corresponding laterality were considered, GPT-4V produced an average PPV of 7.8%, an average TPR of 3.5%, and an average F1 score of 4.5% on the NIH dataset, and an average PPV of 3.6%, an average TPR of 4.9%, and an average F1 of 6.4% on the MIDRC dataset. In few-shot learning, GPT-4V showed improved performance on both datasets. When contrasting zero-shot and few-shot learning, there were improved average TPR and F1 scores in few-shot learning. Nonetheless, there was not a substantial increase in the average PPV.
@@ 2.3 Experimental setup
To obtain the reference standard tables, GPT-4 was used to convert each free-text radiology report (Figure 2a) into a table of radiological findings with the prompt shown in Figure 2b. This table included the radiological findings, the corresponding ICD-10 diagnostic codes and their laterality, and detailed descriptions of the ICD codes (Figure 2c). Subsequently, each report was independently evaluated by three readers from a cohort of five board-certified radiologists and residents (H.O., P.K., C.W., J.K., G.S). Two of the readers were third-year radiology residents, and the remaining three are radiology attendings, each with over 15 years of experience. Their radiology subspecialties cover chest, ED/bone, neurology, and body imaging. All readers had access to the image views and reports but not to additional clinical or patient data. Both datasets, the National Institutes of Health (NIH) and the Medical Imaging and Data Resource Center (MIDRC), were comprehensively reviewed, with each dataset’s 50 reports being examined by three reviewers to maintain consistency and objectivity in the evaluation process. Findings were only included in the final tables if they were observed by at least two of the three reviewers; conversely, findings were excluded if two or more reviewers did not identify them. The majority vote principle was employed to provide a clear consensus for the presence or absence of radiological findings and lead to the final reference-standard table.
@@ 2.3.1 Evaluating Performance in the zero-shot setting
To evaluate GPT-4V’s performance in the zero-shot setting, the chest radiographs (Figure 2d) were input into the GPT-4V model (version from Oct.13th)[2] along with the prompt (Figure 2e). This step aims to generate a radiological findings table (Figure 2f) comparable to the reference standard. The analysis concentrated on aligning positive radiological finding identification and laterality between GPT-generated and consensus tables. Positive predictive value (PPV), true positive rate (TPR), and F1 scores were calculated at the report level (detailed in the section of “Evaluation metrics”). Notably, the application of the F1 scores represented an unconventional approach due to the potential involvement of multiple diagnoses. These metrics were used to assess the accuracy of GPT-4V in detecting the ICD-10 codes and their respective laterality.
@@ 2.3.2 Evaluating Performance in the few-shot setting
To evaluate GPT-4V’s performance in the few-shot setting, the input was extended to include two examples of chest radiographs with their corresponding radiological findings tables before the prompt. From the pool of 10 chest radiographs, two were randomly selected to serve as few-shot examples for the GPT-4V model. Supplying the model with these examples helped to provide context, boosting the model’s capacity to generate an accurate radiological findings table. The same performance metrics— PPV, TPR, and F1 score— were used at the report level to assess the effectiveness of few-shot learning.
@@ 2.4 Statistical Analysis
First, the GPT-4V was employed to detect radiological findings from each chest radiography and compared the predicted findings with those obtained from the radiologists. To convey the performance evaluation, positive predictive value (PPV) was utilized to denote the proportion of ICD-10 Codes correctly predicted by GPT-4V, while the true positive rate (TPR) represented the ration of true positive predictions to the total number of ICD-10 Codes identified by GPT-4V.
Evaluation metrics of PPV, TPR, and F1 score were used to assess GPT-4V’s performance in detecting imaging findings from each chest radiograph [6]. These predicted findings were then compared with those obtained from the radiologists.
@@
For example, in Figure 2, while evaluating ICD-10 codes alone, there were 2 ICD-10 Codes correctly predicted by GPT-4V (J90 and J18.9), five findings predicted by GPT-4V, and four in the reference standard table. Therefore, PPV was 0.4 (=2/(2+3)) , TPR was 0.5 (=2/(2+2)), and F1 score was 0.44.
After obtaining the PPV, TPR, and F1 score for each chest radiography, the macro averages were calculated for PPV, TPR, and F1 score across all 90 chest radiographs.
@@ 3.1 Performance in the zero-shot setting
The performance of the GPT-4V model in the zero-shot setting varied across the NIH and MIDRC datasets and under different circumstances (Figure 3). ). In the task of detecting ICD-10 codes alone, the model attained an average PPV of 5.53/45 (12.3%), with a standard deviation (SD) of 0.25, a standard error (SE) of 0.04, and an interquartile range (IQR) of 0.20; an average TPR of 2.60/45 (5.8%), with a SD of 0.10, a SE of 0.02, and an IQR of 0.10; and an average F1 score of 3.30/45 (7.3%), with a SD of 0.13, a SE of 0.02, and an IQR of 0.14, on the NIH dataset. Conversely, on the MIDRC dataset, the model managed an average PPV of 11.25/45 (25.0%), with a SD of 0.24, a SE of 0.04, and an IQR of 0.33; an average TPR of 7.56/45 (16.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.25; and an average F1 score of 8.20/45 (18.2%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.29. The notable differences in measurements between the two datasets were primarily due to the MIDRC dataset having fewer missing GPT-4V-generated ICD-10 codes than the NIH dataset. On the MIDRC dataset, GPT-4V generated 144 radiological findings, while the reference standard comprised 261 findings. However, on the NIH dataset, GPT-4V produced 102 radiological findings, while the reference standard comprised 220 findings. Nevertheless, when both the ICD-10 codes and their corresponding laterality were taken into account, the GPT-4V model in the zero-shot setting produced an average PPV of 3.5/45 (7.8%), with a SD of 0.20, a SE of 0.03, and an IQR of 0.0; an average TPR of 1.56/45 (3.5%), with a SD of 0.01, a SE of 0.01, and an IQR of 0.0; and an average F1 score of 2.05/45 (4.5%) with a SD of 0.10, a SE of 0.02, and an IQR of 0.0 on the NIH dataset; and an average PPV of 4.92/45 (3.6%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.25; an average TPR of 2.19/45 (4.9%), with a SD of 0.09, a SE of 0.01, and an IQR of 0.10; and an average F1 of 2.90/45 (6.4%), with a SD of 0.11, a SE of 0.02, and an IQR of 0.15, on the MIDRC dataset.
Figure 3: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the zero-shot setting, with statistical significance assessed using a two-tailed t-test. a, Detecting the radiological findings in ICD codes. b, Detecting both the radiological findings in ICD codes and their corresponding lateralities. ICD= International Classification of Diseases, Tenth Revision.
@@ 3.2 Performance in the few-shot setting
In few-shot learning, GPT-4V showed improved performance on both NIH and MIDRC datasets (Figure 4). When the model was provided with two illustrative chest radiographs and reports, there was a marked improvement in the average PPV on the NIH dataset, with rates rising to 5.72/45 (12.7%) with a SD of 0.19, a SE of 0.03, and an IQR of 0.25. Average TPR also enhanced to 4.69/45 (10.4%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22; and the average F1 score reached 5.03/45 (11.1%), with a SD of 0.17, a SE of 0.03, and an IQR of 0.22. On the MIDRC dataset, the average PPV was improved to 16.15/45 (35.9%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; the average TPR improved to 16.68/45 (37.1%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50; and the average F1 score improved to 15.47/45 (34.3%), with a SD of 0.23, a SE of 0.03, and an IQR of 0.50. When tasked with detecting both ICD-10 codes and their corresponding lateralities, the GPT-4V model demonstrated improved efficacy. On the NIH dataset, it displayed an average PPV of 1.62/45(3.5%), with a SD of 0.10, a SE of 0.01, and an IQR of 0.0; an average TPR of 1.14/45 (2.5%), with a SD of 0.07, a SE of 0.01, and an IQR of 0.0; and an average F1 score of 1.30/45(2.8%), with a SD of 0.08, a SE of 0.01, and an IQR of 0.0. On the MIDRC dataset, it achieved an average PPV of 8.96/45 (19.9%), with a SD of 0.22, a SE of 0.03, and an IQR of 0.33; an average TPR of 9.14/45 (20.3%), with a SD of 0.25, a SE of 0.04, and an IQR of 0.33; and an average F1 score of 8.53/45 (19.0%), with a SD of 0.21, a SE of 0.03, and an IQR of 0.31.
Figure 4: Performance of GPT-4 with Vision (GPT-4V) in the detection of radiological findings from chest radiography in the few-shot setting, with statistical significance assessed using two-tailed t-test. a, Detecting the radiological findings in ICD codes. b, Detecting both the radiological findings in ICD codes and their corresponding lateralities. ICD= International Classification of Diseases, Tenth Revision.
When contrasting zero-shot and few-shot learning approaches on both datasets (Figure 5), there were improved average TPR and F1 scores in few-shot learning in both scenarios (ICD-10 codes only and ICD-10 codes with laterality). Nonetheless, there was not a substantial increase in the average PPV suggesting that while few-shot learning may enhance the model’s capacity to detect findings, it does not noticeably enhance the precision of the predictions.
@@ 4 Discussion
The emergence of multimodal large language models (LLMs) that can understand both text and images, such as GPT-4 with Vision (GPT-4V), shows potential for automated image-text pair generation. However, applying these models to real-world data is yet to be thoroughly examined. This study assessed the feasibility of using GPT-4V to detect radiological findings from chest radiographs in both zero-shot and few-shot learning contexts. The results (average PPV of 5.53/45 (12.3%) and average TPR of 2.60/45 (5.8%)) demonstrated that radiological findings tables generated by the GPT-4V model still need to be prepared for use in clinical practice. We acknowledged a limitation in employing GPT-4 for converting radiological reports into a structured table, where inaccuracies in ICD-10 code assignments and distinguishing between radiological findings and conclusions may impact the data’s reliability and interpretability. A notable limitation of the GPT-4V output was its failure to detect several radiological conditions based on corresponding codes in the International Classification of Diseases, Tenth Revision (ICD-10). Overall, the top three findings that GPT-4V could not detect were “endotracheal tube”, “central venous catheter”, and “degenerative changes of osseous structures”. Conversely, GPT-4V most accurately detected findings such as “chest drain”, “air-space disease”, and “lung opacity”.
@@
Acknowledgment
This work was supported by the National Institutes of Health under Award No. 4R00LM013001 (Peng), NSF CAREER Award No. 2145640 (Peng), and Amazon Research Award (Peng, Shih).
We acknowledge parts of this article were generated with GPT-4V (powered by OpenAI’s language model; https://openai.com).